OpenStudy (redprince):

What is the integral of ∫(cos x)^2.sqrt{sin x}dx

OpenStudy (lgbasallote):

\((\cos x)^2 = \cos^2 x = 1 - \sin ^2 x\) i think this can be applied? but seems no :/

OpenStudy (redprince):

@lgbasallote It is not working. I already checked it.....!!

OpenStudy (anonymous):